A sample of gas occupies 600ml at 27 degree Celsius and 1atm. Pressure. What will be the volume at 127 degree Celsius if pressure is Kept constant

Grade:9

1 Answers

Arun
6 years ago
If pressure is constant
Then using charles law
V1/T1 = V2/T2
here Temp in kelvin.
Thus
600/300 = V2/ 400
V2 = 800 ml